OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -6(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which equation best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x)

OpenStudy (anonymous):

Each time x changes by +1, by how much does g(x) change? If this is constant then you have the slope, m, for g(x) = mx + b, The y-intercept, b, is defined as the value for x=0, which you have also from the table. For f(x), the y-intercept is the value f(x=0)= 6(1.02)(0) = 0.

OpenStudy (anonymous):

gx changes plus 2

OpenStudy (anonymous):

so 2=mx + b now what